# Exclusive vector - toponium photoproduction in hadronic collisions

### v1

@@ Abstract
The exclusive production of a vector - toponium $\psi_{t}$ state by photon - hadron interactions is investigated considering proton - proton and proton - nucleus collisions at the Large Hadron Collider (LHC) and Future Circular Collider (FCC) energies. The scattering amplitude is calculated using the $k_{T}$ - factorization formalism assuming that the vector - toponium state can be described by a Gaussian light-cone wave function and considering different models for the unintegrated gluon distribution. Predictions for the rapidity distributions and total cross - sections are presented.
@@ I Introduction
$\frac{d\sigma\,\left[h_{1}h_{2}\rightarrow h_{1}+\psi_{t}+h_{2}\right]}{d^{2}\mbox{\boldmath$b$}\,dy_{\psi_{t}}}=\omega_{h_{1}}N_{h_{1}}(\omega_{h_{1}},b)\,\sigma_{\gamma h_{2}\rightarrow\psi_{t}\otimes h_{2}}\left(\omega_{h_{1}}\right)+(1\longleftrightarrow 2)\,\,$, (1)
@@ II Formalism
In what follows, we will discuss in more details the ingredients needed to calculate the differential cross - section for the exclusive $\psi_{t}$ photoproduction in $h_{1}h_{2}$ collisions, as given in Eq. (1). For the photon spectrum, $N(\omega,b)$, we will consider that it can be expressed in terms in terms of the charge form factor $F(q)$ as follows upc
@@ III Results
In what follows, we will present our predictions for the exclusive $\psi_{t}$ photoproduction in $pp$ and $pPb$ collisions at the LHC and FCC energies. In our calculations, as in Ref. Bai:2025buy , we will assume $m_{t}=172.4$ GeV and $M_{\psi_{t}}=342.1$ GeV. The value of the slope parameter $B$ for the exclusive $\psi_{t}$ photoproduction is an open question. The experimental results from HERA indicate that this value decreases for heavier quarkonium states Ivanov:2004ax . Considering this trending, we expect that $B\leq 4.0$ GeV-2. As this parameter determines the normalization of the cross - section, we will consider, for simplicity, that $B=1$ GeV-2 in Eq. (3), which easily allows the rescaling of our predictions if other value is assumed.
@@
Finally, in Table 1, we present our predictions for the total cross - sections (in ab) associated with the exclusive $\psi_{t}$ photoproduction in $pp$ and $pPb$ collisions at the LHC and FCC energies. We consider the typical rapidity ranges covered by central ($2.5\leq Y\leq+2.5$) and forward ($2.0\leq Y\leq 4.5$) detectors. For $pp$ collisions at the LHC energy, we predict cross - sections of the order of few ab, with the HERAI+II providing a lower bound and being smaller if the rapidity range covered by a forward detector is considered. Such values become of the order of dozen of ab for FCC energies. On the other hand, for $pPb$ collisions at LHC (FCC), we predict an increasing of the cross - section by one (two) order of magnitude in comparison with the $pp$ results. Considering the expected luminosities in the HL-LHC run and in the future hadronic colliders, such results strongly motivates a more detailed analysis about the feasibility of separate these events.
@@ IV Summary
Although the formation of a toponium state in $e^{-}e^{+}$ collisions had been predicted many years ago, the discovery of the top quark with a large mass and rapid decay have reduced the interest about the associated bound state during the last three decades. The situation has recently changed with the observation in $pp$ collisions at the LHC of a substantial excess of events near the kinematical $t\bar{t}$ threshold, which are consistent with the formation of a pseudoscalar $\eta_{t}$ toponium state. Such events have been observed in inclusive processes, where both incident protons fragment in the collision. Considering the large background present in these processes, in Ref. Francener:2025tor the authors have proposed the analysis of the $\eta_{t}$ production in exclusive hadronic interactions, where the incident hadrons remain intact and two rapidity gaps are present in the final state, providing a cleaner environment to separate the toponium state. In this paper, we expand the analysis of exclusive processes as a way to searching for toponium and perform an exploratory study of the exclusive photoproduction of a vector toponium $\psi_{t}$ state. Differently from the $\eta_{t}$ state, which can be exclusively produced by photon - photon and gluon - gluon interactions, the vector state can only be produced by photon - Pomeron interactions. In this paper, we have considered the $k_{T}$ - factorization formalism to describe the exclusive $\psi_{t}$ photoproduction and presented predictions for $pp$ and $pPb$ collisions at the LHC and FCC energies, derived assuming a Gaussian ansatz for the light - cone wave function of meson and considering different parameterizations for the unintegrated gluon distribution. In particular, we presented results for the rapidity distributions and estimates for the total cross - sections considering the typical rapidity ranges covered by central and forward detectors. Our results indicated that the cross - sections are of the order of few ab at the LHC, increasing to some dozen of ab at the FCC. Considering the expected luminosities at the HL - LHC run and at the FCC, our results strongly motivate a more detailed analysis of the process, taking into account of e.g. realistic experimental cuts. Such a study is currently being performed, and the results are expected to be released soon.

### v2

@@ Abstract
An exploratory study of the exclusive production of a vector - toponium $\psi_{t}$ state by photon - hadron interactions is performed considering proton - proton and proton - nucleus collisions at the Large Hadron Collider (LHC) and Future Circular Collider (FCC) energies. The scattering amplitude is calculated using the $k_{T}$ - factorization formalism assuming that the vector - toponium state can be described by a Gaussian light-cone wave function and considering different models for the unintegrated gluon distribution. Predictions for the rapidity distributions and total cross - sections are presented. Our results indicate that the experimental measurement of this final state will be very difficult for the expected integrated luminosities at the LHC and FCC.
@@ I Introduction
$\frac{d\sigma\,\left[h_{1}h_{2}\rightarrow h_{1}+\psi_{t}+h_{2}\right]}{d^{2}\mbox{$b$}\,dy_{\psi_{t}}}=\omega_{h_{1}}N_{h_{1}}(\omega_{h_{1}},b)\,\sigma_{\gamma h_{2}\rightarrow\psi_{t}\otimes h_{2}}\left(\omega_{h_{1}}\right)+(1\longleftrightarrow 2)\,\,$, (1)
@@ II Formalism
In what follows, we will discuss in more details the ingredients needed to calculate the differential cross - section for the exclusive $\psi_{t}$ photoproduction in $h_{1}h_{2}$ collisions, as given in Eq. (1). For the photon spectrum, $N(\omega,b)$, we will consider that it can be expressed in terms of the charge form factor $F(q)$ as follows upc
@@ III Results
In what follows, we will present our predictions for the exclusive $\psi_{t}$ photoproduction in $pp$ and $pPb$ collisions at the LHC and FCC energies. In our calculations, as in Ref. Bai:2025buy , we will assume $m_{t}=172.4$ GeV and $M_{\psi_{t}}=342.1$ GeV. The value of the slope parameter $B$ for the exclusive $\psi_{t}$ photoproduction is an open question. The experimental results from HERA indicate that this value decreases for heavier quarkonium states Ivanov:2004ax . Considering this trending, we expect that $B\leq 4.0$ GeV-2. As this parameter determines the normalization of the cross - section, we will consider, for simplicity, that $B=1$ GeV-2 in Eq. (3), which easily allows the rescaling of our predictions if other value is assumed. However, it is important to emphasize that such assumption for the value of $B$ implies that our predictions for the magnitude of the cross - section are optimistic and must be considered as an upper bound.
@@
Finally, in Table 1, we present our predictions for the total cross - sections (in ab) associated with the exclusive $\psi_{t}$ photoproduction in $pp$ and $pPb$ collisions at the LHC and FCC energies. We consider the typical rapidity ranges covered by central ($2.5\leq Y\leq+2.5$) and forward ($2.0\leq Y\leq 4.5$) detectors. For $pp$ collisions at the LHC energy, we predict cross - sections of the order of few ab, with the HERAI+II providing a lower bound and being smaller if the rapidity range covered by a forward detector is considered. Such values become of the order of dozen of ab for FCC energies. On the other hand, for $pPb$ collisions at LHC (FCC), we predict an increasing of the cross - section by one (two) order of magnitude in comparison with the $pp$ results.
It is important to emphasize that the experimental separation of the events associated with the exclusive quarkonium photoproduction is, in general, performed using the dileptons generated in the quarkonium decay. Considering the expected luminosities in the HL-LHC run and in the future hadronic colliders and the results derived in Ref. (Bai:2025buy, ) for the branching fraction of the vector toponium into dileptons ${\cal{O}}(10^{-5})$], the predicted number of events at the LHC and FCC will be strongly suppressed, implying that the experimental measurement of this final state will be, in principle, a very hard task. Currently, we are investigating alternatives to separate these events and intend to present the results in a forthcoming study.
@@ IV Summary
Although the formation of a toponium state in $e^{-}e^{+}$ collisions had been predicted many years ago, the discovery of the top quark with a large mass and rapid decay have reduced the interest about the associated bound state during the last three decades. The situation has recently changed with the observation in $pp$ collisions at the LHC of a substantial excess of events near the kinematical $t\bar{t}$ threshold, which are consistent with the formation of a pseudoscalar $\eta_{t}$ toponium state. Such events have been observed in inclusive processes, where both incident protons fragment in the collision. Considering the large background present in these processes, in Ref. Francener:2025tor the authors have proposed the analysis of the $\eta_{t}$ production in exclusive hadronic interactions, where the incident hadrons remain intact and two rapidity gaps are present in the final state, providing a cleaner environment to separate the toponium state. In this paper, we expand the analysis of exclusive processes as a way to searching for toponium and perform an exploratory study of the exclusive photoproduction of a vector toponium $\psi_{t}$ state. Differently from the $\eta_{t}$ state, which can be exclusively produced by photon - photon and gluon - gluon interactions, the vector state can only be produced by photon - Pomeron interactions. In this paper, we have considered the $k_{T}$ - factorization formalism to describe the exclusive $\psi_{t}$ photoproduction and presented predictions for $pp$ and $pPb$ collisions at the LHC and FCC energies, derived assuming a Gaussian ansatz for the light - cone wave function of meson and considering different parameterizations for the unintegrated gluon distribution. In particular, we presented results for the rapidity distributions and estimates for the total cross - sections considering the typical rapidity ranges covered by central and forward detectors. Our results indicated that the cross - sections are of the order of few ab at the LHC, increasing to some dozen of ab at the FCC. Considering the expected luminosities at the HL - LHC run and at the FCC and the quarkonium decay into dileptons, our results indicated that a future experimental separation of this final state will be, in principle, a very hard task. Such result motivates a more detailed analysis of the process, considering new strategies to separate the associated events as well as taking into account of e.g. realistic experimental cuts. Such a study is currently being performed, and the results are expected to be released soon.